### v2

Abstract
Speculative decoding in large language models (LLMs) accelerates token generation by speculatively predicting multiple tokens cheaply and verifying them in parallel, and has been widely deployed. In this paper, we provide the first study demonstrating the privacy risks of speculative decoding. We observe that input-dependent patterns of correct and incorrect predictions can be leaked out to an adversary monitoring token generation times and packet sizes, leading to privacy breaches. By observing the pattern of correctly and incorrectly speculated tokens, we show that a malicious adversary can fingerprint queries and learn private user inputs with more than 90% accuracy across three different speculative decoding techniques – REST (almost 100% accuracy), LADE (up to 92% accuracy), and BiLD (up to 95% accuracy). We show that an adversary can also leak out confidential intellectual property used to design these techniques, such as data from data-stores used for prediction (in REST) at a rate of more than 25 tokens per second, or even hyper-parameters used for prediction (in LADE). We also discuss mitigation strategies, such as aggregating tokens across multiple iterations and padding packets with additional bytes, to avoid such privacy or confidentiality breaches.
1 Introduction
Large Language Models (LLMs) have transformed natural language processing (NLP), enabling machines to generate and understand human language at an unprecedented scale Vaswani et al. (2017); Devlin et al. (2019); Brown et al. (2020); Zhang et al. (2022); Le Scao et al. (2022); Touvron et al. (2023). LLMs typically generate text using auto-regressive decoding Touvron et al. (2023); OpenAI (2024a), where the generation happens serially and each token depends on all the previous ones. Unfortunately, this serial process causes a significant bottleneck in the LLM response latency Miao et al. (2024); Fu et al. (2024) and under-utilizes the available hardware-level parallelism, limiting token generation throughput and latency.
Speculative decoding Leviathan et al. (2023); Chen et al. (2023); Miao et al. (2024); Spector & Re (2023) addresses this problem without impacting model accuracy. It uses smaller models or heuristics such as retrieval or self-drafting to inexpensively generate tokens speculatively, which the larger target model verifies in parallel in a single iteration. By tuning the heuristics to maintain a high rate of correct speculations, such techniques provide 2x to 5x speedups in inference latency and throughput Xia et al. (2024).
However, the adoption of speculative techniques is not without risks. Speculative execution in CPUs  Burton (1985); Hennessy & Patterson (2012), which inspired speculative decoding, has led to considerable security vulnerabilities in processors, such as Spectre Kocher et al. (2019), Meltdown Lipp et al. (2018), and their numerous variants. These side-channel vulnerabilities exploit the side-effects of processor mis-speculations, such as timing variations introduced by mis-speculations, to leak private and confidential data that become accessible to a processor during mis-speculations. This raises the question, does speculative decoding in LLMs also introduce new risks to privacy?
In this paper, we provide the first study of the privacy and confidentiality risks of speculative decoding in LLMs, including leakage of private user inputs or confidential data.
Problem. We observe two key issues in speculative decoding implementations: (1) the pattern of correct and incorrect speculations of output tokens is dependent on the input, and (2) input-dependent speculation patterns can be inferred from the variations in packet sizes: upon correct speculation, more tokens generated per iteration leads to a larger packet and so observing the packet size leaks the degree of correct speculation. Consequently, an adversary capable of measuring the number of tokens generated per iteration or packet sizes can gain access to input-dependent speculation-patterns based on variation in packet sizes and leak out private input attributes, or even entire inputs and outputs.
Q. How to fine-tune a large language model?   A.  F ine - t uning _a _large _language _model _involves _adjust ing _the _parameters _of _the _model _to _improve _its _performance _on _a _specific _task _or _down stream _task . _Here _are _some _steps _you _can _follow _to _fine - t une _a _large _language _model : <0x0A> <0x0A> 1 . _Cho ose _the _down stream _task : _Once _you _have _a _large _language _model , ...
(a) Speculative Decoding Example
(b) Tokens Per Autoregressive Iteration
(c) Potential Exploits
Figure 1: (a) In an LLM response with Speculative Decoding (e.g., LADE), tokens are either correctly speculated (blue) and verified in parallel, or mis-speculated (red) and generated via auto-regressive decoding. (b) This pattern can be inferred by measuring the number of tokens generated per iteration, where multiple tokens per iteration indicates correct speculation, and a single token per iteration indicates mis-speculation. (c) Using these patterns, a network-based adversary can fingerprint user queries and learn private user prompts and responses; a malicious user can observe correct predictions and leak out data-stores and hyper-parameters used for predictions.
Consider a response generated by TinyLlama 1.1B Chat V1.0 Zhang et al. (2024a) with the state-of-the-art speculative decoding technique, Lookahead Decoding (LADE) Fu et al. (2024) in Figure 1(a), as an example. The red tokens are mis-speculated, requiring one auto-regressive iteration per token. Whereas, the blue tokens are correctly speculated, with up to $n$ tokens verified in a single iteration, as shown in Figure 1(b). Thus, in LLM applications that stream responses back to the user iteration by iteration, an adversary over the network, similar to recent attacks on LLMs Weiss et al. (2024), can identify network packets corresponding to each iteration and observe the packet size to deduce the number of tokens produced by each iteration and leak the pattern of correct and incorrect speculations dependent on the input.
Exploits. By observing speculation patterns, we demonstrate three possible exploits on speculative-decoding implementations, as shown in Figure 1(c).
First, we demonstrate a query fingerprinting attack by a network-based adversary. By using the speculation-pattern as a fingerprint for an prompt/response pair, the adversary can profile the fingerprints for different prompt/output pairs offline, and then for an unknown user query, measure the speculation-pattern, and compare it with the profiled fingerprints to leak out the user’s private prompt and response. On three different speculative decoding implementations (BiLD, REST, LADE) Fu et al. (2024); He et al. (2024); Kim et al. (2023), we show that one can identify user inputs using fingerprints with accuracy of up to $\sim$100for REST, 92% for LADE, and 95% for BiLD with just 30 training samples. When the prompts to be leaked are just approximately similar to profiled prompts, the accuracy is still within 20% to 40% accuracy (much better than random guessing probability of 2%).
Second, we show that an adversary capable of crafting malicious inputs can leak confidential data or hyper-parameters that may be used for generating predictions in speculative decoding techniques. If predictions are generated via retrieval from a data-store (REST) He et al. (2024), we show that an adversary can craft inputs and monitor which output tokens are correctly predicted, and leak out the data-store at the rate of >25 tokens per second. When speculative decoding generates predictions using self-drafting (LADE) Fu et al. (2024), we show that an adversary can craft malicious inputs that leak prediction hyperparameters like the N-gram size, and the number of prediction candidates generated and verified, which are confidential intellectual property controlling the success rates of a speculative-decoding implementation.
Mitigations. Finally, we discuss potential mitigations against these exploits. We propose transmitting groups of tokens aggregated over several iterations back to the user instead of streaming tokens for each iteration to reduce the observability of fine-grain speculation patterns, and padding packets with a fixed number or random number of tokens to add noise to adversary’s observations to mitigate the fingerprinting attacks. Avoiding usage of private or confidential data for generating predictions also prevents leakage.
In summary, our contributions are as follows:
  1. 1.
    We are the first to observe that variation in the number of tokens generated per autoregressive iteration due to LLM speculative decoding can result in privacy and confidentiality breaches.
  2. 2.
    We demonstrate query fingerprinting attacks that can leak out exact matches of private user queries with >90% accuracy and approximate matches with 20% to 40% accuracy.
  3. 3.
    We also demonstrate attacks that leak confidential IP controlling the performance of the speculative decoding mechanisms, such as data from data-stores used for prediction or hyper-parameters used for predictions.
  4. 4.
    We also propose mitigations including padding packet sizes and aggregating tokens from multiple iterations to limit the potential for these exploits.
2 Background and Motivation
We provide background on speculative decoding in LLMs, and then cover recent side-channel attacks on LLMs.
2.1 Background on Speculative Decoding
Speculative decoding addresses inefficiencies in auto-regressive decoding by allowing multiple token predictions to be verified in parallel, reducing overall latency. It works by generating speculative token guesses, which are then verified by the main model. If correct, multiple tokens are accepted in one step, significantly speeding up the process. If incorrect, the model falls back to standard autoregressive decoding to ensure accuracy.
Speculation Using Smaller Models. Many works propose using smaller, much faster draft models to generate sequences of tokens speculatively Miao et al. (2024); Kim et al. (2023), and then verify a single sequence or a tree of such sequences in parallel using the larger target model. The draft model can be a smaller version of the model family or a pruned larger model Yan et al. (2024). In this paper, we use BiLD Kim et al. (2023), as a representative example of this type. Two key policies—fallback and rollback—ensure the quality of tokens is maintained in BiLD: the fallback policy hands control to the larger model if confidence drops, while rollback corrects tokens exceeding a set distance metric, like cross-entropy loss.
Speculation using Self-Drafting. Another speculative decoding paradigm caches previous output tokens and retrieves them for speculation in future iterations, with the target model verifying these speculations Cai et al. (2024); Fu et al. (2024). Look-Ahead Decoding (LADE) Fu et al. (2024) implements this by maintaining a cache where keys are tokens and values are N-grams previously generated after those tokens. LADE populates its cache using jacobi decoding, selecting random tokens from the input and generating token chains. When a key token reappears in the output stream, LADE uses these N-grams for speculation.
Speculation Using Retrieval. Unlike other paradigms relying on model generations, Retrieval-Based Speculative Decoding (REST) He et al. (2024) speculates by retrieving data from a datastore containing contextually relevant text or code. REST’s datastore contains prefixes and continuations from a corpus, matching input suffixes to generate draft outputs. If there are too many drafts, REST selects high-quality subsets and organizes them in a tree, applying a tree attention mask during decoding with the target model.
2.2 Recent Side-Channel Attacks on LLMs
Recent works have demonstrated side-channel vulnerabilities in LLM inference systems leading to critical privacy risks. A recent attack Weiss et al. (2024) demonstrated a token-length side-channel on streaming LLMs, where network-based attackers can leak a large fraction of private response tokens from an LLM chatbot, by observing the size of encrypted packets in the network traffic. By identifying individual tokens based on packet sizes and leveraging inter-sentence context, this attack reconstructs a significant portion of private responses from platforms like ChatGPT and Microsoft Copilot that stream responses token-by-token. Another attack  Debenedetti et al. (2024) demonstrated a tokenizer side-channel, where a malicious user constructs inputs that translate to different number of tokens based on the tokenizer’s vocabulary, which is used to leak out the vocabulary and any private data in the vocabulary.
2.3 Goal: Speculative Decoding Side-Channel
Speculative execution techniques in processors, deployed for improving performance, have resulted in many potent side-channel attacks capable of leaking information, such as Spectre and Meltdown Kocher et al. (2019); Lipp et al. (2018) and their many variants. Consequently, this work seeks to explore whether the deployment of speculative decoding in LLMs similarly for improving inference performance introduces any new side-channels capable of privacy breaches.
3 Threat Model
Like prior works Weiss et al. (2024), we focus on streaming LLMs, where the user sends requests to an LLM running on the server, and the server sends back the response token-by-token. Here, we study the following attacker models.
  1. 1.
    Network Based Adversary. We assume the attacker is able to observe the network packet size and the time at which they are transmitted, similar to Weiss et al. (2024). The goal of the attacker is to leak private user queries. Although network packets containing LLM response tokens are encrypted, the packet size is still observable. The attacker can also observe the time between each packet to identify which iteration it belongs to. We show this allows an adversary to fingerprint and leak private user queries in Section 4.
  2. 2.
    Malicious User. In this setting, we assume the attacker is a malicious user interacting with the LLM, similar to Debenedetti et al. (2024), with the goal of leaking confidential data used by LLM’s speculative decoding mechanism. Here, the user can craft arbitrary inputs and inspect the output, and also monitor the number of tokens generated per iteration for the output tokens. We show this allows an adversary to leak out any confidential data used for the speculative decoding or its hyperparameters in Section 5.
4 Query Fingerprinting Attack
4.1 Attack Overview
Attack Scenario. We consider a scenario where a user interacts with an AI chatbot using an LLM. The LLM is programmed to answer a predefined set of user queries through the system prompt. This can occur in scenarios such as a patient in a healthcare setting interacting with an LLM to get medical queries answered, or a user in a customer service setting sharing private information with the chatbot before being redirected to a human. The queries are sent from a client device, over a network, to a server that hosts the LLM equipped with speculative decoding. This leads to the threat of a network-based attacker trying to eavesdrop on the queries and responses, as described in Section 3.
Figure 2: Overview of the Query Fingerprinting Attack. In the offline phase, the network-based attacker profiles the trace of number of tokens generated per iteration influenced by speculative decoding, for all possible inputs, and trains a classifier. In the online phase, the attacker extracts the trace of number of tokens generated per iteration for an unknown input, and uses the trained classifier to leak the input.
Goals and Capabilities. Figure 2 shows the overview of our attack. Our attacker over the network seeks to learn the private query asked by the user. While the query and response packets are encrypted, the attacker can intercept the response packets, measure the time intervals between packets, and count the number of tokens per packet or simply use the packet size to approximate the number of tokens.
Mechanism. Since the input set is limited, the attacker fingerprints inputs based on the pattern of output token counts per iteration influenced by speculative decoding. There are two phases to the attack: offline and online. In the offline phase, the attacker profiles the number of output tokens generated per iteration vs iteration-id (we call this a trace) for each input in the input set, and trains a random forest classifier to predict the input based on the trace. In the online phase, the attacker collects the trace for an unknown input and uses the classifier to identify the input. Next, we provide more details about how these fingerprints are obtained, and then describe how the attack works on various speculative decoding techniques (LADE, REST, BiLD).
4.2 Fingerprinting Using Speculation Patterns
Speculative decoding introduces variation in the number of tokens generated per iteration. In iterations with correct speculation, multiple tokens are verified in a single iteration and sent back together in a single packet. However, if the speculation is incorrect, a single token is generated per iteration. The attacker uses this variation to create a fingerprint for a particular output sequence.
As shown in  Figure 2, from each packet containing outputs from a single iteration, the attacker measures the number of tokens generated per iteration. Subsequently, by plotting the token numbers vs iteration ID, it gets the trace for an input/output (prompt-response) tuple.
Figure 3 shows the trace of tokens per iteration for two different prompts with LADE across different runs. From this, we observe two key properties that make these traces suitable for usage to identify and leak the prompt and response:
  1. 1.
    Traces are unique to a prompt and response - different prompts result in different traces of tokens per iteration, as shown in Figure 3(a) and Figure 3(b).
  2. 2.
    Traces are reproducible - repeating the same prompt produces similar patterns in the trace of token counts, as shown in the two iterations in Figure 3(a) or the two iterations in Figure 3(b).
(a) What is a narcissistic personality disorder?
(b) What are the symptoms of cancer?
Figure 3: Trace of tokens generated per iteration vs Iteration-ID, for two runs of different prompts (a) and (b) with LADE. The trace has unique and reproducible patterns, enabling its usage as a fingerprint for leaking prompts.
These observations demonstrate that the number of tokens generated per iteration with speculative decoding is a unique identifier for a prompt/response pair. These patterns can leak the prompt/response via side-channels to an adversary.
However, in existing LLM deployments on servers such as OpenAI’s ChatGPT, the logits get decoded into character strings by the detokenization at the server-side before being streamed to the user in encrypted packets. Because the packets are encrypted, the tokens in each packet cannot be directly inferred. Nevertheless, there is a strong correlation between the packet size and the number of tokens per iteration. Figure 4 shows the trace of token count per iteration for a prompt with LADE, and the corresponding packet size in bytes per iteration match considerably. Thus, for our fingerprinting attack, the adversary can use the trace of packet size per iteration as a proxy for the token count per iteration, given that it is easily observable.
Figure 4: Comparing the trace of token count per iteration and trace of packet size, versus iteration-id for a prompt response with LADE. Strong similarity between the traces allows the attacker to infer speculation patterns, i.e., variation in tokens per iteration, via variation in packet size.
Next, we develop attacks to fingerprint and leak user inputs, based on the traces of packet sizes per iteration.
4.3 Attack Experiment Design
We perform our attacks on an AI Chatbot designed to answer common medical queries from a patient. The user is a patient interacting with the chatbot and the attacker seeks to learn which query the user asks. We conduct the following three fingerprinting attacks:
Experiment 1. The first experiment studies a scenario where the attacker has exact knowledge of the complete set of queries, allowing the attacker to profile the exact queries offline that will be leaked out in the online phase. In this, we use a set of 50 prompts from real world human-LLM interactions acquired from Han et al. (2023). These prompts ask a variety of questions about different common diseases (the list of prompts is provided in the Appendix).
Experiment 2. In this experiment, we choose 50 prompts with commonplace diseases from our dataset that all starts with “What are the symptoms of”. Therefore, all the prompts have a similar structure. We design this experiment to see if our attack can distinguish similar prompts.
The training and testing prompts of both Experiment 1 and 2 are the same. This illustrates the case where the attacker has exact knowledge of the set of predefined prompts.
Experiment 3. The last experiment studies a scenario where the attacker has only approximate knowledge of the set of queries asked by the user. To emulate this, the attacker profiles queries that are semantically similar to the queries asked by the user (e.g., profiled query: “For how long should you take zolpidem?” and user: “How long is zolpidem typically prescribed for?”). We reuse the prompts from Experiment 1 as our training prompts, but for testing, i.e., the user prompts to be leaked, we rephrase the prompts in the training set using ChatGPT-4o by prompting it as "Can you rephrase the following 50 questions while keeping the diseases in each question the same?"
4.4 Methodology
Target Models. We target three recent speculative decoding techniques: LADE Fu et al. (2024) that uses self-drafting for speculation, REST He et al. (2024) that uses retrieval from datastore for speculation, and BiLD Kim et al. (2023) which uses a smaller model for speculation. However, our results are as applicable to other techniques using speculation. We demonstrate results for LADE with TinyLlama 1.1B Chat V1.0 Zhang et al. (2024a), for REST with Vicuna 7B Chiang et al. (2023) as the target model and ShareGPT Aeala (2023) data-store containing 120K conversations used for speculation, and for BiLD with Llama2 7B Touvron et al. (2023) as the target model and TinyLlama 1.1B Zhang et al. (2024a) as the smaller model for speculation. We run LADE and REST on NVIDIA RTX4090 (24GB) and BiLD on NVIDIA A100 (40GB) GPUs with Question-Answer tasks, to simulate a chatbot under attack.
Datasets and Classifier. For each experiment, we use 50 queries of varying complexity, length, and semantics (Appendix lists all the prompts). For our training set, in the offline profiling phase of the attack, we run each query 5 to 30 times and extract the packet size traces (5 to 30 traces per query) and train a Random Forest classifier using them. For the test set in the online phase, we use the corresponding 50 queries and 5 new traces per query, collecting 250 new data points for evaluation, and report accuracy and F1-scores. For our attack, we use scikit-learn’s Random Forest with 150 estimators, max depth of 15, minimum 10 samples per split, minimum 1 sample per leaf, and mean squared error as the loss function.
Accuracy (%) F1 Score
TPQ $\rightarrow$ 5 10 20 30 5 10 20 30
Experiment 1
LADE 53.6% 63.2% 65.5% 68% 0.49 0.59 0.63 0.66
REST 91.2% 97.6% 100% 99.6% 0.91 0.98 1.0 1.0
BiLD 28.4% 41.6% 49.2% 52% 0.27 0.43 0.52 0.55
Experiment 2
LADE 33.2% 53.2% 64.8% 70.8% 0.3 0.51 0.63 0.7
REST 96.4% 97.6% 100% 99.6% 0.96 0.98 1.0 1.0
BiLD 58.8% 68% 75.2% 77.6% 0.54 0.64 0.73 0.76
Experiment 3
LADE 15.2% 16% 13.6% 18.8% 0.13 0.13 0.1 0.16
REST 35.2% 38.4% 40% 39.6% 0.27 0.32 0.34 0.33
BiLD 13.2 % 20% 20.4% 20% 0.11 0.16 0.16 0.16
Table 1: Results of the Query Fingerprinting Attack on three speculative decoding techniques (LADE, REST, BiLD) with temperature of 0.8, with 5 to 30 Traces Per Query (TPQ) used for training the classifier.
4.5 Results
Table 1 shows the results for the Query Fingerprinting Attack on LADE, REST, and BiLD.
Exact Knowledge Scenarios (EK). In both Experiment 1 and 2, the attacker leaks the input prompt with high accuracy: with accuracy of up to 70.8%, 100%, and 77.6% for LADE, REST, and BiLD respectively, compared to random guessing (2% for a set of 50 prompts). In Experiment 1, where we study randomly selected prompts and the set of profiled and leaked prompts are identical, the attack success rates improve as the number of traces per query (TPQ) profiled increases, reaching up to 68% accuracy and an F1 score of 0.66 with 30 TPQ for LADE, reaching 99.6% accuracy with 30 TPQ and an F1 score of 1.0 for REST, and reaching 52% accuracy with 30 TPQ and an F1 score of 0.55 for BiLD. In Experiment 2, where all our prompts have a similar structure, the results are similar, with attack accuracy for LADE, REST, and BiLD reaching 70.8%, 99.6%, and 77.6% respectively. This indicates that our attack success rate is not influenced by the structure of the prompt (input), as long as the response (output) is different.
Approximate Knowledge Scenarios (AK). In Experiment 3, where the leaked prompts are not exactly the same as the profiled prompts, the attack success rates are lower (up to 18.8% for LADE, 20% BiLD, and 40% for REST), but still significantly better than random-guessing (2%). This demonstrates that models using speculative decoding are vulnerable to query fingerprinting attacks, even when the attacker just has approximate knowledge about the queries.
Overall, all of the attacks are more successful on REST than BiLD and LADE. This is because REST demonstrates higher correct speculation rates compared to BiLD and LADE, and therefore has more prompt-dependent variations in the tokens per iteration and the packet sizes, leading to the attack achieving higher accuracy on REST.
4.6 Ablation Studies
Table  2 presents an ablation study studying the attack accuracy as the temperature of our target model varies.
Temperature
0.3 0.6 0.8 1.0
Experiment 1
LADE 91.6% 85.6% 68% 61.2%
REST 100% 100% 99.6% 99.6%
BiLD 74.4% 59.2% 52% 34.8%
Experiment 2
LADE 90.8% 84.4% 70.8% 60.4%
REST 99.6% 100% 99.6% 100%
BiLD 95.2% 88.4% 77.6% 63.6%
Experiment 3
LADE 16.8% 17.6% 18.8% 19.6%
REST 35.2% 42% 39.6% 43.2%
BiLD 17.2% 15.6% 20% 16.4%
Table 2: Accuracy of the Query Fingerprinting Attack on LADE and REST at different temperatures (0.3, 0.6, 0.8, 1.0), with 30 Traces Per Query (TPQ) used to train the classifier.
For LADE and BiLD, as the temperature increases from 0.3 to 1.0, the attack accuracy for both Experiment 1 and Experiment 2 decreases. This is because at lower temperatures, the generations and speculation patterns are more stable and alike. Therefore, the output tokens for the same prompt and the packet size traces are more similar across different runs, making the attack more successful.
For REST, the attack accuracy is relatively stable for both Experiment 1 and 2 despite the increasing temperature. This is likely because the features are so pronounced that it withstands the perturbations caused in the output by the increasing temperature.
The attack accuracy in Experiment 3 has an interesting trend with temperature. Since the attacker only has approximate knowledge of the prompts (train and test set are different), the accuracy of this attack is much lower than Experiment 1 and 2. However, unlike Experiment 1 and 2, for Experiment 3 the attack accuracy has a moderate increase as the temperature of the model increases. This is because, at low temperatures, with limited variation in the training data, the classifier overfits on the training set, and at higher temperatures, this problem does not exist, allowing the attack to reach slightly higher accuracies.
5 Intellectual Property Leakage
Speculative decoding techniques may use carefully selected data or tuned hyper-parameters for achieving the highest correct speculation rates. In this section, we focus on how a malicious user can craft inputs to recover such data or hyper-parameters. We assume the malicious user can examine the output and count the number of tokens streamed back per iteration, as outlined in the threat model in Section 3-2.
5.1 Leaking Datastore Used for Speculation
Techniques like REST He et al. (2024) rely on retrieval from an external datastore to generate speculative tokens. These datastores can be populated with proprietary information to tailor model outputs to specific domains, or data collected from user interactions that is private. Patterns of correct speculations arising from such data can be exploited to leak the sensitive information in such datastores.
5.1.1 Background on REST’s Speculation
REST builds a datastore of phrases in the format of context and continuation pairs, where a preceding context and its following continuation are stored. During generation, REST retrieves the longest context matching with the tokens generated so far, and uses it to construct a speculative draft by selecting top prefixes from the retrieved continuations. This speculative draft is verified using a tree attention mechanism by the target model.
5.1.2 Attack Design
We observe that any tokens that are correctly speculated, i.e., returned as a group of tokens per iteration have to exist in the datastore. We propose an attack where a malicious user crafts inputs and observes which tokens are correctly speculated and all such tokens can thus be leaked from the datastore. With this, the attacker can systematically extract sequences from the datastore.
We study the following attacker strategies for input generation to achieve high token leakage rates:
1. Random Generation: The user prompts the LLM to generate random paragraphs of text.
2. Leveraging Common Words. The LLM is prompted to generate paragraphs containing the most frequently occurring words, taken from a dataset of top 10,000 English words Brants & Franz (2021). This is based on the insight that there is a high chance of phrases in the datastore containing these words.
3. Reusing Leaked Sequences. The LLM is prompted to generate paragraphs that extend phrases leaked out so far, as shown in Figure 5. This approach is based on the insight that the sequence that is already leaked out may be a part of a longer phrase in the datastore. Therefore, with such iterative prompting, additional tokens from the phrase can be leaked out in a focused manner, leading to more unique tokens being leaked with a similar number of queries.
Figure 5: Leakage process where an attacker exploits leaked sequences to extract longer datastore sequences.
Figure 6: Unique sequences leaked from REST datastore as time increases with different attack strategies, averaged over 3 runs per strategy.
5.1.3 Results
We test our attacks on a REST implementation using the ShareGPT datastore of 120K conversations for speculation on NVIDIA RTX 4090. Figure 6 shows the unique sequences leaded by each of the three attack strategies, averaged over three runs per strategy.
Random generation leaks 20,000 unique sequences in 20 minutes, but subsequently the rate slows down leaking only 70,000 sequences in 3 hours. In comparison, prompting with frequently occurring words leaks 31,000 unique sequences in 20 minutes, which increases linearly to 190,000 after 3 hours. Lastly, prompting with leaked sequences results in the most efficient leakage, leaking about 35,000 unique sequences in 20 minutes and about 200,000 after 3 hours, highlighting the amplification potential of feedback. These results show that an attacker can easily leak REST’s datastore when they can observe correct speculation patterns.
5.2 Leakage of Speculation Hyper-parameters
Speculative decoding mechanisms rely on hyper-parameters to control speculation, directly impacting correct speculation rates and performance. By analyzing traces of these patterns, we can reverse-engineer the hyper-parameters used. We demonstrate this attack on LADE Fu et al. (2024).
5.2.1 Background on LADE’s Implementation
LADE generates speculative tokens by caching and reusing prior n-grams generated by the model. LADE has two main parameters: N for n-gram size, and G for guess set size. Its cache is structured as a key-value store, where the key is a token, and the value is a list of G candidate (N-1)-grams following the key token, that are used for speculation. These are replaced in the Least-Recently-Used (LRU) order. During execution, LADE use the last token of the previous iteration or the input to query this cache, and the associated G candidate (N-1)-grams are used as speculation candidates, and the best match with target model generation is accepted.
5.2.2 Leaking Hyper-parameter N in LADE
As each iteration outputs at most N-1 tokens under correct speculation, we expect the maximum number of tokens per iteration to be N-1. Therefore, we prompt LADE with a simple prompt, “Repeat Letter ‘A’ 60 times” that should sustain the maximum correctly speculated tokens per iteration. We observe the maximum number of tokens per iteration with this prompt to deduce N-1, and thus learn N.
Figure 7 shows the number of correct speculated tokens for LADE with three different values of $N$ for this prompt. We see that for each value of N (4, 5, and 6), this prompt has a maximum of 3, 4, and 5 tokens per iteration (N-1), letting us learn the value of N as 4, 5, or 6 in each of these cases. This technique can be extended to learn any value of N used in LADE.
Figure 7: Number of tokens per iteration with LADE for different N, with a prompt that achieves maximum correct speculation (“Repeat Letter ‘A’ 60 times”. The maximum token count per iteration corresponds to value of N-1.)
5.2.3 Leaking Hyper-parameter G in LADE
To leak G, the number of candidate predictions considered for speculation, we force LADE to incur targeted mis-speculations. Here, we prompt the model to repeat a sequence of phrases with the same prefix (“I run”), i.e., “I run to the grocery shop. I run as fast as cars. I run with my best friend. I run from the lecture hall.". If the number of such phrases is less than or equal to G, LADE caches all the N-grams following the key “run”, resulting in it being able to correctly speculate the token that follows “run” in each of these phrases. However, if the number of phrases are greater than G, then with LRU replacement, newer n-grams evict older n-grams for the key “run” from the set of candidates which has a limited capacity of G. This causes the token following “run” to always be mis-speculated.
Figure 8 shows the traces of token generation time for LADE ($G=3$), as the number of phrases ($P$) of the type “I run …” increases from 2 to 4. When $P$ is 2 or 3 ($\leq G$), we notice all the tokens after the initial ones are correctly predicted in the generation (high number of tokens per iteration). But when $P$ is 4 ($>G$), every 7th token, i.e. the token following “run” is mis-speculated (one token per iteration). Thus, we leak the value of G to be 3. This attack can be generalized to any value of parameter G in LADE.
Figure 8: Time Per Token vs Token-ID for LADE with Guess Set (G = 3), using a prompt that asks the LLM to repeat P phrases sharing a common prefix “I run”. Periodic mis-speculations occur for the token after “run”, only when $P>G$, leaking the value of $G$.
6 Mitigation Strategies
In this section, we discuss possible mitigation strategies for our attacks and evaluate their impact on the attack success rate as appropriate.
6.1 Mitigating Query Fingerprinting Attacks
For the query fingerprinting attacks, we propose the following mitigations:
1. Padding Token Packets: Padding the network packet size with additional bytes can mask the actual number of tokens per iteration and conceal the differences between correct and mis-speculations to prevent the attack.
The padding can be added in two ways:
  1. 1.
    Constant Length Padding. Every packet can be padded to a constant size. This size can be selected based on the maximum number of tokens per iteration, multiplied by the maximum bytes per token. We pick this size as 1024 bytes, without loss of generality.
  2. 2.
    Variable Length Padding. Every packet can be padded with a random number of bytes selected from a uniform distribution of $\epsilon\sim Unif(0,D)$ where D is a pre-defined upper bound of padding and $\epsilon$ is an integer.
Table 3: Accuracy after padding each packet to 1024 bytes (attack: 20 traces per prompt; model temperature: 0.8)
No Padding Padding to 1024 Bytes
Experiment 1
LADE 65.6% 3.2%
REST 100% 2%
BiLD 49.2% 2%
Experiment 2
LADE 64.8% 2.4%
REST 100% 2%
BiLD 75.2% 2%
Experiment 3
LADE 13.6% 4%
REST 40% 2%
BiLD 20.4% 2%
Constant Length Padding. Table 3 shows the accuracy of the attack after padding every response packet to a constant size of 1024 Bytes. This completely mitigates the side channel, reducing the attack accuracy to close to random guessing (2%). However, this comes at the cost of considerable communication overhead and inefficiency, due to an increase in packet size by almost 230$\times$, as majority of the packet consists of random padding bytes.
Table 4: Attack accuracy after padding each packets with random bytes from uniform distribution, $\epsilon\sim Unif(0,D)$, where D varies from 6 to 48. (attack: 20 traces per prompt; model temperature: 0.8)
Accuracy vs Max Padding Size (D) Overhead vs D
D $\rightarrow$ 0 6 12 24 48 6 12 24 48
Experiment 1
LADE 65.6% 25.6% 13.6% 8.4% 3.6% 1.5$\times$ 2$\times$ 3.1$\times$ 5.1$\times$
REST 100% 90.4% 75.6% 48% 27.2% 1.9$\times$ 2.8$\times$ 4.6$\times$ 8.2$\times$
BiLD 49.2% 26.8% 11.6% 5.2% 4.8% 1.7$\times$ 2.3$\times$ 3.6$\times$ 6.3$\times$
Experiment 2
LADE 64.8% 19.2% 10.8% 5.6% 4.4% 1.6$\times$ 2.1$\times$ 3.2$\times$ 5.4$\times$
REST 100% 95.2% 76.8% 52% 34.4% 2$\times$ 3$\times$ 4.8$\times$ 8.7$\times$
BiLD 75.2% 37.2% 21.6% 8.4% 5.6% 1.7$\times$ 2.4$\times$ 3.7$\times$ 6.5$\times$
Experiment 3
LADE 13.6% 6.8% 6% 3.6% 3.2% 1.5$\times$ 2$\times$ 3$\times$ 5$\times$
REST 40% 29.6% 20.8% 14.8% 8.4% 1.9$\times$ 2.8$\times$ 4.6$\times$ 8.1$\times$
BiLD 20.4% 7.6% 6.8% 6% 3.6% 1.7$\times$ 2.3$\times$ 3.6$\times$ 6.2$\times$
Variable Length Padding. Table 4 shows the attack accuracy after padding responses with a variable number of bytes, $\epsilon\sim Unif(0,D)$, where D varies from 6 to 48. As D increases to 48, the attack accuracy for LADE and BiLD drops close to random guess (2%) for all experiments. The attack accuracy for REST also shows a considerable drop from 100% without padding to 27% and 34% with padding of up to 48 bytes for Experiments 1 and 2 respectively. Similarly, for Experiment 3, attack accuracy drops from 40% to 8% with D=48. These reductions in accuracy come with a 5$\times$ to 8$\times$ increase in the packet size for D=48. Thus, variable length padding provides a better trade-off of reduction in attack accuracy with limited increase in packet size compared to constant length padding.
2. Reduce Granularity of Attacker Observation: An alternative approach is to aggregate output tokens over multiple iterations at the server and return them to the client all together. This can limit observability of fine-grain speculation patterns from the packet sizes for the attacker, and reduce the effectiveness of fingerprinting attacks.
Table 5 shows the results for the server aggregating output tokens over 3 to 20 iterations. The accuracy of the exact knowledge attacks (Experiment 1 and 2) for LADE, REST, and BiLD drops to a minimum of 9.2%, 52%, and 11.2% respectively after aggregating the tokens over up to 20 iterations down from 65%, 100%, and 49.2% respectively without aggregation; accuracy of the approximate knowledge attacks (Experiment 3) decreases to a minimum of 5.2%, 17.6%, and 4% respectively, down from 14%, 40%, and 20% respectively.
Thus, reducing granularity is an effective mitigation against query fingerprinting attacks. One downside of this approach is that the user waits longer between updates to the output as the number of iterations the tokens are aggregated over increases. This limits the number of iterations tokens can be aggregated over without hindering system responsiveness.
Table 5: Accuracy of fingerprinting attacks (20 traces per prompt and temperature of 0.8) with aggregation of tokens across 3 to 20 iterations.
Aggregation Granularity (Number of Iterations)
1 3 5 10 20
Experiment 1
LADE 65.6% 53.2% 40% 24.4% 10.8%
REST 100% 96.8% 89.2% 80.4% 52%
BiLD 49.2% 34% 30% 19.2% 11.2%
Experiment 2
LADE 64.8% 45.2% 28.8% 16.4% 9.2%
REST 100% 97.6% 92.8% 75.6% 52%
BiLD 75.2% 57.6% 41.6% 30% 16.8%
Experiment 3
LADE 13.6% 13.2% 14.4% 9.2% 5.2%
REST 40% 33.2% 28.8% 20.4% 17.6%
BiLD 20.4% 15.6% 10% 10% 4%
6.2 Mitigating Datastore or Hyper-parameter Leaks
For preventing confidential data leakage from datastore or parameter leakage attacks, in addition to the two mitigations discussed previously, we can consider one other mitigation:
3. Using Only Public Data for Speculative Decoding: Any data-store used for speculation (e.g., in REST) should be constructed only with public data, and any private data or personal identifiable information (e.g., usernames, address, etc.) should be scrubbed from it before it is used. This minimizes the risk of sensitive data leakage through any speculation patterns that may be observable despite padding or token aggregation. Similarly, parameters used for speculation (such as N and G in LADE) should be assumed to be public as they can be easily leaked out.
7 Related Work
In this section, we describe other speculative decoding techniques that may be exploitable and related LLM performance optimizations that may lead to side-channel leakage.
7.1 LLM Speculative Decoding Techniques
Speculative Decoding with Tree Attention. SpecInfer Miao et al. (2024) and SpecTr Sun et al. (2023) observe that when a token early in the speculation draft is rejected, the entire draft gets discarded, resulting in wasted computation. Instead, these works propose multiple drafts per iteration organized in a tree structure to provide alternatives for speculative tokens at each position, increasing the likelihood of tokens advancing to the next step and yielding greater speedup by increasing correct speculation rates. These drafts are organized in a tree structure with a custom tree attention mask to prevent inter-tree interactions and verified in parallel. Such works (similar to REST He et al. (2024) which also uses tree attention) increase the correct speculation rate of tokens and introduce more input-dependent speculation patterns in the output. Hence these are likely to be more vulnerable to our query fingerprinting attacks, similar to REST, which is highly susceptible to our attack, as we show in this paper.
Speculative Decoding using Target Models. Other works reuse the target model itself for speculative decoding by utilizing the feature layer (second-to-top layer). Medusa Cai et al. (2024) trains multiple extra heads, each responsible for a token position in the draft sequence, to generate speculative tokens. EAGLE and EAGLE-2 Li et al. (2024a; b) propose a method to perform auto-regressive decoding on the feature layers itself to generate speculative tokens. These techniques are conceptually no different from using an alternative smaller model for speculation like BiLD. Hence, they are equally susceptible to query fingerprinting attacks.
7.2 Performance Optimizations in LLMs
Other performance optimization techniques in LLMs that can introduce timing variation, may also be susceptible to leakage via timing side-channels.
Adaptive KV-Cache Management. The KV-cache is the bottleneck for token generation latency for the memory-bound LLM inference Pope et al. (2023). Recent works, like H2O Zhang et al. (2024b) that performs input-dependent cache eviction or TriForceSun et al. (2024) that performs retrieval-based KV-cache selection for drafting may have input-dependent timing variations. These can introduce timing side-channels allowing query fingerprinting, similar to our attacks. However, as such attacks rely on timing, they are likely to be more susceptible to noise.
Prompt Caching A significant portion of prompts sent to LLMs are duplicates, including system prompts, repeated inputs, and others. Prompt caching, adopted by major LLM providers such as Anthropic and OpenAI Anthropic (2024); OpenAI (2024b); Gim et al. (2024), precomputes attention states for these repetitive prompt components to enable reuse, and substantially reducing latency for the first token when there is reuse. Such techniques can leak the nature of the query based on the timing. Other works cache the responses for prompts with similar semantic meanings Bang (2023). Recent attacks Song et al. (2024) demonstrate query fingerprinting attacks similar to ours, when such prompt caches are shared between users. In contrast, our attacks are effective even without any explicit sharing of caches.
8 Conclusion
This paper reveals significant privacy risks associated with speculative decoding in Large Language Models. Across multiple speculative decoding techniques, we demonstrate that attacks can exploit speculation patterns to infer user inputs with $>90$% accuracy via fingerprinting attacks and also leak data from datastores used for predictions. These leakages highlight the careful consideration needed when deploying speculative decoding techniques to ensure that performance does not come at the cost of reduced privacy.
Appendix A Prompts for Fingerprinting Attacks
A.1 Training and testing prompts for Experiment 1
What to expect if I have Allergy (Outlook/Prognosis)?
What causes Cardiogenic shock?
What to expect if I have Congenital heart block (Outlook/Prognosis)?
What are the symptoms of Congestive heart failure?
What are the symptoms of Coronary heart disease?
What to expect if I have Dengue fever (Outlook/Prognosis)?
What causes Stroke?
Who is at highest risk for Heart attack?
What to expect if I have Heart murmur (Outlook/Prognosis)?
When to seek urgent medical care when I have Metabolic syndrome?
What are the symptoms of Palpitation?
What causes Radiation injury?
What are the symptoms of shock?
Who is at highest risk for Back pain?
What are the symptoms of Athlete’s foot?
What causes Boil?
When to seek urgent medical care when I have Burns?
What causes Corns & calluses?
Who is at highest risk for Fleas?
When to seek urgent medical care when I have Hand-foot-and-mouth disease?
When to seek urgent medical care when I have Hives?
When to seek urgent medical care when I have Liver spots?
What to expect if I have Measles (Outlook/Prognosis)?
When to seek urgent medical care when I have Mumps?
What are the symptoms of a Rash?
When to seek urgent medical care when I have Ringworm?
What are the symptoms of Scarlet fever?
What causes Shingles?
What are the symptoms of Sunburn?
When to seek urgent medical care when I have Wart?
What are the symptoms of Common cold?
Who is at highest risk for AIDS ?
What are the causes of Alcoholic liver disease?
What to expect if I have Anal cancer (Outlook/Prognosis)?
When to seek urgent medical care when I have Anal fissure?
When to seek urgent medical care when I have Anemia?
When to seek urgent medical care when I have Antisocial personality disorder?
How many children have Autism?
What are the symptoms of Avian influenza?
When to seek urgent medical care when I have Avoidant personality disorder?
What causes cataract?
Who is at highest risk for Chickenpox?
What causes Chronic fatigue syndrome?
When to seek urgent medical care when I have Depression?
What to expect if I have Color blindness?
How many children have Dependent personality disorder?
When to seek urgent medical care when I have Adolescent depression?
Who is at highest risk for Diabetes?
What to expect if I have Ebola?
What causes fatty liver?
A.2 Training and testing prompts for Experiment 2
What are the symptoms of Allergy?
What are the symptoms of Cardiogenic shock?
What are the symptoms of heart block?
What are the symptoms of Congestive heart failure?
What are the symptoms of Coronary heart disease?
What are the symptoms of Dengue fever?
What are the symptoms of stroke?
What are the symptoms of a Heart attack?
What are the symptoms of Heart murmur?
What are the symptoms of Metabolic syndrome?
What are the symptoms of Palpitation?
What are the symptoms of Radiation injury?
What are the symptoms of shock?
What are the symptoms of Back pain?
What are the symptoms of Athlete’s foot?
What are the symptoms of Boil?
What are the symptoms of Burns?
What are the symptoms of Corns & calluses?
What are the symptoms of Fleas?
What are the symptoms of Hand-foot-and-mouth disease?
What are the symptoms of Hives?
What are the symptoms of Liver spots?
What are the symptoms of Measles?
What are the symptoms of Mumps?
What are the symptoms of a Rash?
What are the symptoms of Ringworm?
What are the symptoms of Scarlet fever?
What are the symptoms of Shingles?
What are the symptoms of Sunburn?
What are the symptoms of Warts?
What are the symptoms of Common cold?
What are the symptoms of AIDS?
What are the symptoms of Alcoholic liver disease?
What are the symptoms of Anal cancer?
What are the symptoms of an Anal fissure?
What are the symptoms of Anemia?
What are the symptoms of Antisocial personality disorder?
What are the symptoms of Autism?
What are the symptoms of Avian influenza?
What are the symptoms of Avoidant personality disorder?
What are the symptoms of a cataract?
What are the symptoms of Chickenpox?
What are the symptoms of Chronic fatigue syndrome?
What are the symptoms of Depression?
What are the symptoms of Color blindness?
What are the symptoms of Dependent personality disorder?
What are the symptoms of Adolescent depression?
What are the symptoms of Diabetes?
What are the symptoms of Ebola?
What are the symptoms of fatty liver?
A.3 Training prompts for Experiment 3
Same as Section A.1
A.4 Testing prompts for Experiment 3
What is the outlook for someone with an allergy?
What leads to cardiogenic shock?
How might congenital heart block affect me in the long term?
How does congestive heart failure present itself?
How can coronary heart disease symptoms be identified?
What is the prognosis for someone with dengue fever?
What are the main causes of a stroke?
Who faces the greatest risk for a heart attack?
What is the long-term prognosis for someone with a heart murmur?
When is it critical to get help for metabolic syndrome?
How can I tell if I’m experiencing palpitations?
What factors contribute to radiation injury?
What are the warning signs of shock?
Who is most susceptible to back pain?
What should I look out for if I suspect athlete’s foot?
What are the main causes of a boil?
When are burns severe enough to seek emergency medical care?
Why do corns and calluses form?
Who is most at risk of getting fleas?
When should I go to the doctor for hand-foot-and-mouth disease?
When is hives a sign to seek urgent medical attention?
When should liver spots be evaluated by a doctor?
What can I expect if I have measles?
When should I seek medical help for mumps?
How do I know if I have a rash?
When should ringworm be treated urgently?
What are the common symptoms of scarlet fever?
What is the underlying cause of shingles?
How can I recognize sunburn symptoms?
When should a wart be looked at by a doctor?
What are the signs of the common cold?
Who is most at risk of contracting HIV/AIDS?
What factors contribute to alcoholic liver disease?
What is the outlook for someone diagnosed with anal cancer?
When is an anal fissure an emergency?
When should anemia symptoms prompt immediate medical care?
When is antisocial personality disorder considered an urgent issue?
How common is autism in children?
What symptoms are typical of avian influenza?
When should avoidant personality disorder be addressed by a professional?
What are the main causes of cataracts?
Who is most vulnerable to contracting chickenpox?
What are the underlying causes of chronic fatigue syndrome?
When should I seek emergency medical help if I have depression?
What are the typical symptoms or outcomes if I have color blindness?
How common is dependent personality disorder in children?
When is urgent medical attention needed for adolescent depression?
Who faces the highest risk of developing diabetes?
What should I expect if I have Ebola?
What leads to the development of fatty liver?